OpenStudy (anonymous):

how do I solve for x lim5 if X to 3rd minus 125 over x-5 if i am solving for the limit of the function

ganeshie8 (ganeshie8):

\[\large \lim \limits_{x \to 5} ~\dfrac{x^3-125}{x-5}\]

ganeshie8 (ganeshie8):

like that ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

il give a hint : \(\large a^3 - b^3\) formula

ganeshie8 (ganeshie8):

another hint : 125 = 5^3